OpenStudy (anonymous):

simplify to the lowest terms 16a^2/4a+11b-121b^2/4a+11b

OpenStudy (anonymous):

\[\frac{ 16a^2 }{ 4a + 11 } - \frac{ 121b^2 }{ 4a + 11b }\] i'm guessing you meant that?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

find an LCD which in this case is?

OpenStudy (anonymous):

i don't know...nothing goes into 121

OpenStudy (anonymous):

the LCD is least common denominator, so the left fraction's denominator is the same as the right 4a + 11b is the LCD

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

so how do I find the lowest term ?

OpenStudy (anonymous):

is the answer 105/4a+11b ?

OpenStudy (anonymous):

since the right side's denominator is already 4a + 11b we have to multiply the left sides to look like that as well we're missing a B so we multiply the numerator and denominator by B/B so it now looks like \[\frac{ 16a^2b }{ 4a + 11b } - \frac{ 121b^2 }{ 4a + 11b }\]

OpenStudy (anonymous):

\[\frac{ 16a^2 - 121b^2 }{ 4a + 11b }\]

OpenStudy (anonymous):

ok